OpenStudy (anonymous):

Turner’s treadmill starts with a velocity of −2.7 m/s and speeds up at regular intervals during a half-hour workout.After 33 min, the treadmill has a velocity of −7.1 m/s. What is the average acceleration of the treadmill during this period? Answer in units of m/s2

OpenStudy (anonymous):

It starts with a negative velocity?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You can't have a negative velocity.

OpenStudy (anonymous):

If you're moving, then your velocity is always positive.

OpenStudy (anonymous):

The lowest your velocity can ever be if measured from an inertial referent is 0

OpenStudy (anonymous):

velocity is a vector so it can be negative or positive

OpenStudy (anonymous):

Not true. Depends on your inertial reference frame and axis of choice...

OpenStudy (anonymous):

Okay, the paper says it has negative velocity.

OpenStudy (anonymous):

It only makes sense geometrically, not physically

OpenStudy (anonymous):

Can anybody help me on this?

OpenStudy (anonymous):

in any case the answer is: -(7.1-2.1)/(33*60)

OpenStudy (anonymous):

just take the difference in speed. then divide it by the ammount of passed seconds.

OpenStudy (anonymous):

use the newtons equations for constant acceleration v = u + a t -7.1 = -2.7 + a * 33 * 60 - sove for a

OpenStudy (anonymous):

i.e. if you have a speed, [m/s] and you have a time interval [min]=[60*s] and you need an answer in [m/s^2]

OpenStudy (anonymous):

Does anyone have the answer?